OpenStudy (anonymous):

When a car hits a certain bump on the road, a shock absorber on the car is compressed a distance of 8 in., then released. The shock absorber vibrates in damped harmonic motion with a frequency of 6 cycles per second. The damping constant for this particular shock absorber is 3.4. I found the equation of f(t)=8e^-3.4tcos(12πt). How long does it take for the amplitude of the vibration to decrease to 0.8 in.? (Round your answer to two decimal places.)

OpenStudy (aum):

Assuming the equation for f(t) is correct, the amplitude is \(\large 8e^{-3.4t}\). Equate the amplitude to 0.8 and solve for t.

OpenStudy (anonymous):

The equation for f(t) is correct but I'm stuck on trying to solve for t. My professor never went over the chapter that this is in.

OpenStudy (aum):

\[ 8e^{-3.4t} = 0.8 \\ \text{Divide by 8:} \\ e^{-3.4t} = 0.1 \\ \text{Take ln on both sides:} \\ -3.4t = \ln(0.1) \\ t = \frac{\ln(0.1)}{-3.4} = ? \]

OpenStudy (anonymous):

Ah, thank you!